# A faster algorithm for Vertex Cover parameterized by solution size

### v6

@@ 10.4 Line 8: Branching on a 5-vertex with two 6-neighbors
In the algorithm Branch6-2, we consider a 5-vertex $u$ with 6-neighbors $y_{1},y_{2}$ and 5-neighbors $v_{1},v_{2},v_{3}$, where either $y_{1}$ or $y_{2}$ has at most three 5-neighbors. See Figure 9 for an illustration.
@@
  • Vertices $v_{1},v_{2},v_{3}$ are independent, as otherwise $R(y_{1},y_{2})\geq 1$ (due to 3-triangle $u$) and we would branch at Line 3. There must be some pair $y_{i},v_{j}$ with $y_{i}\not\sim v_{j}$ (else both $y_{1},y_{2}$ would have four 5-neighbors $u,v_{1},v_{2},v_{3}$); say without loss of generality that $y_{1}\not\sim v_{1}$. We have $y_{2}\not\sim\{v_{1},v_{2},v_{3}\}$, else $R(y_{1},v_{1})\geq 1$ due to the 3-triangle involving $u,y_{2},v_{j}$, and we would apply Line 3 to $y_{1},v_{1}$. This in turn implies that $y_{1}\not\sim\{v_{2},v_{3}\}$, else $R(y_{2},v_{1})\geq 1$ due to the resulting 3-triangle and we would apply Line 3 to $y_{2},v_{1}$.
  • Say without loss of generality that $y_{1},v_{1}$ share a vertex $x\neq u$. Since none of the neighbors of $u$ are adjacent to $v_{1}$, we have $x\not\sim u$. If $\deg(x)\leq 4$, then $R(y_{1},v_{1})\geq 1$ (due to subquadratic vertex $x$), and we would branch on $y_{1},v_{1}$ at Line 3. If $\deg(x)=6$ and $x\sim y_{2}$, then $R(N[x])\geq 1$ (due to vertex $u$ losing neighbors $y_{1},v_{1},y_{2}$) and we would branch at Line 2. If $\deg(x)=6$ and $x\not\sim y_{2}$, then $R(y_{2},N[x])\geq 1$ (again due to vertex $u$ losing neighbors $y_{1},v_{1},y_{2}$), and we would branch on $y_{2},x$ at Line 7.
  • Suppose $y_{i}$ shares neighbor $t_{ij}\neq u$ with $v_{j}$ for $j=1,2,3$; as we have just seen, $\deg(t_{ij})=5$ and $t_{ij}\not\sim u$ for all $j$. Vertices $t_{i1},t_{i2},t_{i3}$ must be distinct; for, if $t_{ij}=t_{ij^{\prime}}=t$, then $\operatorname{codeg}(t,u)\geq 3$ (sharing vertices $v_{j},v_{j^{\prime}},y_{i}$), contradicting Observation 61. So $y_{i}$ has four distinct 5-neighbors $t_{i1},t_{i2},t_{i3},u$. By hypothesis, this does not hold for both $y_{1}$ and $y_{2}$. ∎
@@
Suppose $y_{1}\sim y_{2}$, and let $i\in\{1,2\},j\in\{1,2,3\}$. For the graph $G^{\prime}_{ij}=G-\{v_{1},v_{2},v_{3}\}-N[y_{i},v_{j}]$, we have $\operatorname{minsurp}(G^{\prime}_{ij})\geq-3+\operatorname{codeg}(y_{i},v_{j})$. Moreover, if $\operatorname{codeg}(y_{i},v_{j})=1$ and $\operatorname{minsurp}(G^{\prime}_{ij})=-2$, then either $S(G^{\prime}_{ij})\geq 1$ or $R(v_{j^{\prime}},N[v_{j}])\geq 1$ for some $j^{\prime}\in\{1,2,3\}$.
@@
Now suppose $\operatorname{codeg}(y_{1},v_{1})=1$ and $\operatorname{surp}_{G^{\prime}}(I)=-2$. If $|I|\geq 3$ then $S(G^{\prime})\geq 1$ as desired, so suppose $|I|=2$. By Observation 61, we have $\operatorname{codeg}(v_{2},v_{3})\leq 2$, so there must be some vertex $x\in I$ and $v_{j}\in\{v_{2},v_{3}\}$ with $x\not\sim v_{j}$; say without loss of generality $x\not\sim v_{3}$ and $x$ is isolated in $G^{\prime}$.
Let $d=\deg_{G}(x)$. By Observation 61 we have $\operatorname{codeg}(x,y_{1})\leq d-3$ and $\operatorname{codeg}(x,v_{1})\leq 2$. The only way for $x$ to be isolated in $G^{\prime}$ is to have $x\sim v_{2},\operatorname{codeg}(x,v_{1})=2,\operatorname{codeg}(x,y_{1})=d-3$. If $d\leq 5$, we thus have $R(v_{2},N[v_{1}])\geq 1$ due to subquadratic vertex $x$. If $d=6$, then by Proposition 66 we have $x\not\sim y_{2}$ (else the 6-vertex $x$ would be shared by $y_{2},v_{2}$); in this case, $G-N[x]$ has a 3-vertex $y_{1}$ with a triangle $y_{1},y_{2},u$, and we would branch on $x$ at Line 2. ∎
@@
Case I: $\boldsymbol{y_{1}\not\sim y_{2}}$. Split on $y_{1},y_{2},v_{1}$, getting subproblems $G_{1}=G-N[y_{1}],G_{2}=G-y_{1}-N[y_{2}],G_{3}=G-\{y_{1},y_{2}\}-N[v_{1}],G_{4}=G-\{y_{1},y_{2},v_{1}\}$. By Observation 61, we have $\operatorname{minsurp}(G_{1})\geq 2$ and $\operatorname{minsurp}(G_{2})\geq 1$ and $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(y_{1},v_{1})-1\geq 1$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(y_{1},y_{2})-1\geq 1$. So $G_{1},G_{2},G_{3},G_{4}$ have drops $(2.5,6),(3,7),(3,7),(1.5,3)$ directly. We also apply Proposition 30 to 2-vertex $u$ in $G_{4}$, where $\deg(v_{2})+\deg(v_{3})-1-\operatorname{codeg}(v_{2},v_{3})\geq 7$ by Observation 61. So we need:
@@
Case II: $\boldsymbol{y_{1}\sim y_{2}}$. We choose some enumeration $v_{1},v_{2},v_{3}$, and split on $v_{1},v_{2}$; in the subproblem $G-\{v_{1},v_{2}\}$, we apply (A2) to $u,y_{1}$. This generates subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G\{v_{1},v_{2}\}-N[v_{3},y_{1}],G_{4}=G-\{v_{1},v_{2},y_{1}\}$.
By Observation 61, $G_{1},G_{2},G_{4}$ have drops $(2,5),(2.5,6),(1.5,3)$ directly; furthermore, we can apply Proposition 30 to the 2-vertex $u$ in $G_{4}$, where $r=\deg(v_{1})+(\deg(y_{2})-1)-\operatorname{codeg}(y_{2},v_{1})-1\geq 5+5-2-1=7$ by Observation 61.
Proposition 66 and Observation 61 give $\operatorname{minsurp}(G_{3})\geq-3+\operatorname{codeg}(y_{1},v_{3})$ and $\operatorname{codeg}(y_{1},v_{3})\leq 2$. In light of Proposition 65, there are two cases to select the ordering:
Case IIa: $\boldsymbol{R(v_{1},N[v_{2}])\geq 1.}$ In this case $G_{2}$ has drop $(2.5,7)$ after simplification. Since $G_{3}$ has drop $(4,11)$, we need:
@@
Case IIb: $\boldsymbol{\operatorname{codeg}(y_{1},v_{3})=1.}$ By Proposition 66, either $R(v_{j},N[v_{3}])\geq 1$ for some $j$ (which would put us into Case IIa, with some other vertex ordering), or $S(G_{3})\geq 1$, or $\operatorname{minsurp}(G_{3})\geq-1$. In the latter two cases, $G_{3}$ has drop $(4,13)$ or $(4.5,12)$. So we need:
@@
Simplify $G$ Branch on a vertex $u$ with degree larger than 7 Split on a 7-vertex $u$ with $\operatorname{shad}(N[u])\geq 0$. Branch on a 7-vertex $u$ which has a blocker $x$ with $\deg(x)\geq 5$. Branch on a vertex $x$ which is a blocker of three 7-vertices $u_{1},u_{2},u_{3}$. Run either Branch6-3 or the MaxIS-7 algorithm.
Algorithm 10 Function Branch7$(G,k)$
@@
Line 2. Here, we apply Section 5; we need:
@@
Line 3. In this case $G-N[u]$ has drop $(3,7)$. So we need:
@@
Line 4. Apply (B) to $u,x$, getting subproblems $G-N[x],G-\{u,x\}$ with drops $(1,5)$ and $(1,2)$ respectively. We need:
@@
Line 5. Apply (B) to vertices $u_{1},u_{2},u_{3},x$. The resulting subproblems $G-\{u_{1},u_{2},u_{3},x\}$ and $G-N[x]$ have drops $(1,4),(1,3)$ directly; the branch-seq $[(1,3),(1,4)]$ is already covered in (7-1).
@@
If $\operatorname{maxdeg}(G)\leq 7$, then use Lemma 10 with the algorithms Branch7 or MaxIS-7. Otherwise, split on a vertex of degree at least $8$.
Algorithm 11 Function Branch8$(G,k)$
@@
Splitting on a vertex of degree $r\geq 8$ generates subproblems with vertex covers of size at most $k-1$ and $k-8$. These have cost $1.25284^{k-1}$ and $1.25284^{k-8}$ by induction. Since $11.25284^{-1}+1.25284^{-8}<$ 1, the overall cost is $O^{*}(1.25284^{k})$. Otherwise, Lemma 10 (with parameters $a_{7},b_{7}$ and $c=\log(1.19698)$) gives the desired algorithm with runtime $O^{*}(1.25284^{k})$. ∎
@@ Appendix A Proofs for basic results
Proof of Proposition 4.
Consider any basic solution $\theta\in\{0,\tfrac{1}{2},1\}^{V}$ to $\operatorname{LPVC}(G)$. Let $I_{0}=\theta^{-1}(0)$ and $I_{1}=\theta^{-1}(1)$; note that $I_{0}$ is an indset (possibly empty) and $N(I_{0})\subseteq I_{1}$. So $\sum_{v}\theta(v)=|I_{1}|+\tfrac{1}{2}(n-|I_{0}|-|I_{1}|)\geq(n+|N(I_{0})|-|I_{0}|)/2=(n+\operatorname{surp}(I_{0}))/2\geq(n+\operatorname{minsurp}^{-}(G))/2$. On the other hand, if $I_{0}$ is a min-set of $G$, then setting $\theta(v)=1$ for $v\in N(I_{0})$ and $\theta(v)=0$ for $v\in I_{0}$ and $\theta(v)=1/2$ for all other vertices $v$ gives a fractional vertex cover with $\sum_{v}\theta(v)=(n+\operatorname{minsurp}^{-}(G))/2$. ∎
@@
Let $I$ be a critical-set; we first claim there is a good cover $C$ with $I\subseteq C$ or $I\cap C=\emptyset$. For, suppose that $I\cap C=J$ where $\emptyset\neq J\subsetneq I$. In this case, we must also have $N(I\setminus J)\subseteq C$, so overall $|C\cap N[I]|\geq|J|+|N(I\setminus J)|=|J|+|I\setminus J|+\operatorname{surp}(I\setminus J)$; since $I\setminus J$ is a non-empty subset of $I$, we have $\operatorname{surp}(I\setminus J)\geq\operatorname{surp}(I)$, hence this is at least $|J|+|I\setminus J|+\operatorname{surp}(I)=|I|+\operatorname{surp}(I)\geq N(I)$. Thus, replacing $C\cap N[I]$ by $N(I)$ would give a vertex cover $C^{\prime}$ with $C^{\prime}\cap I=\emptyset$ and $|C^{\prime}|\leq|C|$.
@@
Proof of Proposition 6.
Let $J$ be an indset with $\operatorname{surp}_{G-N[I]}(J)=\operatorname{minsurp}^{-}(G-N[I])$. Then $J\cup I$ is a non-empty indset of $G$ with surplus $\operatorname{surp}_{G-N[I]}(J)+\operatorname{surp}_{G}(I)=\operatorname{minsurp}^{-}(G-N[I])+\operatorname{surp}_{G}(I)$; in particular, if $\operatorname{minsurp}^{-}(G-N[I])+\operatorname{surp}_{G}(I)=\operatorname{minsurp}(G)$, then $I$ is a min-set of $G$. On ther other hand, if $I$ is contained in a min-set $J$ of $G$, then $\operatorname{surp}_{G-N[I]}(J\setminus I)=\operatorname{surp}_{G}(J)+|I|-|N(I)|=\operatorname{surp}_{G}(J)-\operatorname{surp}_{G}(I)$. So $\operatorname{minsurp}^{-}(G-N[I])\leq\operatorname{minsurp}(G)-\operatorname{surp}_{G}(I)$. ∎
Proof of Proposition 7.
By applying Proposition 6 to indsets $I=\{x\}$, we get $\operatorname{minsurp}(G)=\min_{x\in V}\bigl{(}\operatorname{minsurp}^{-}(G-N[x])+\deg(x)-1\bigr{)}$. We can use Proposition 4 to compute each value $\operatorname{minsurp}^{-}(G-N[x])$ in polynomial time (by solving the LP). This gives an algorithm to compute $\operatorname{minsurp}(G)$. Likewise, given any vertex-set $X$, we can use Proposition 6 to determine if $X$ is contained in a min-set of $G$ and, if so, we can use Proposition 4 to find a vertex set $J$ attaining $\operatorname{surp}_{G-N[X]}(J)=\operatorname{minsurp}^{-}(G-N[X])$; then $J\cup X$ is a min-set of $G$ containing $X$. ∎
@@
Proof of Lemma 9.
Consider a kite $u,x,y,z$. We can view the application of (P3) to funnel $u$ as a two-part process. First, we remove the shared neighbor $y\in N(u)\cap N(x)$; then, we merge vertices $z,x$ into a new vertex $u^{\prime}$. In the first step, we obtain a graph $G^{\prime\prime}=G-y$ with $k^{\prime\prime}=k-1$ and $n^{\prime\prime}=n-1$. So, by Proposition 4, we have $\lambda(G^{\prime\prime})=\lambda(G)-1/2$ and hence $\mu(G^{\prime\prime})\leq\mu(G)-1/2$ and $k^{\prime\prime}=k-1$. By Proposition 8, the second step gives $\mu(G^{\prime})\leq\mu(G^{\prime\prime})$ and $k^{\prime}=k^{\prime\prime}-1$. ∎
Proof of Lemma 10.
First, exhaustively apply rule (P1) to $G$; the resulting graph $G^{\prime}$ has $n^{\prime}\leq n$ and $k^{\prime}\leq k$, and $\operatorname{minsurp}(G^{\prime})\geq 0$. In particular, $\mu(G^{\prime})=k^{\prime}-\lambda(G^{\prime})=k^{\prime}-n^{\prime}/2$. Note that $G^{\prime}\in\mathcal{G}$ since (P1) just deletes vertices. Now dovetail the two algorithms for $G^{\prime}$, running both simultaneously and returning the output of whichever terminates first. This has runtime $O^{*}(\min\{e^{cn^{\prime}},e^{a(k^{\prime}-n^{\prime}/2)+bk^{\prime}}\})$. For fixed $k^{\prime}$, this is maximized at $n^{\prime}=\frac{2k^{\prime}(a+b)}{a+2c}$, at which point it takes on value $e^{dk^{\prime}}\leq e^{dk}$. ∎

### v7

@@ 8.4 Line 8: Branching on a 5-vertex with two 6-neighbors
In the algorithm Branch6-2, we consider a 5-vertex $u$ with 6-neighbors $y_{1},y_{2}$ and 5-neighbors $v_{1},v_{2},v_{3}$, where either $y_{1}$ or $y_{2}$ has at most three 5-neighbors. See Figure 12.
@@
  • Vertices $v_{1},v_{2},v_{3}$ are independent, as otherwise $S(y_{1},y_{2})\geq 1$ (due to 3-triangle $u$) and we would branch at Line 3. There must be some pair $y_{i},v_{j}$ with $y_{i}\not\sim v_{j}$ (else both $y_{1},y_{2}$ would have four 5-neighbors $u,v_{1},v_{2},v_{3}$); say without loss of generality that $y_{1}\not\sim v_{1}$. We have $y_{2}\not\sim\{v_{1},v_{2},v_{3}\}$, else $S(y_{1},v_{1})\geq 1$ due to the 3-triangle involving $u,y_{2},v_{j}$, and we would apply Line 3 to $y_{1},v_{1}$. This in turn implies that $y_{1}\not\sim\{v_{2},v_{3}\}$, else $S(y_{2},v_{1})\geq 1$ due to the resulting 3-triangle and we would apply Line 3 to $y_{2},v_{1}$.
  • Say without loss of generality that $y_{1},v_{1}$ share a vertex $x\neq u$. Since none of the neighbors of $u$ are adjacent to $v_{1}$, we have $x\not\sim u$. If $\deg(x)\leq 4$, then $S(y_{1},v_{1})\geq 1$ (due to subquadratic vertex $x$), and we would branch on $y_{1},v_{1}$ at Line 3. If $\deg(x)=6$ and $x\sim y_{2}$, then $S(N[x])\geq 1$ (due to vertex $u$ losing neighbors $y_{1},v_{1},y_{2}$) and we would branch at Line 2. If $\deg(x)=6$ and $x\not\sim y_{2}$, then $S(y_{2},N[x])\geq 1$ (again due to vertex $u$ losing neighbors $y_{1},v_{1},y_{2}$), and we would branch on $y_{2},x$ at Line 7.
  • Suppose $y_{i}$ shares neighbor $t_{ij}\neq u$ with $v_{j}$ for $j=1,2,3$; as we have just seen, $\deg(t_{ij})=5$ and $t_{ij}\not\sim u$ for all $j$. Vertices $t_{i1},t_{i2},t_{i3}$ must be distinct; for, if $t_{ij}=t_{ij^{\prime}}=t$, then $\operatorname{codeg}(t,u)\geq 3$ (sharing vertices $v_{j},v_{j^{\prime}},y_{i}$), contradicting Observation 62. So $y_{i}$ has four distinct 5-neighbors $t_{i1},t_{i2},t_{i3},u$. By hypothesis, this does not hold for both $y_{1}$ and $y_{2}$. ∎
@@
Suppose $y_{1}\sim y_{2}$, and let $i\in\{1,2\},j\in\{1,2,3\}$. For the graph $G^{\prime}_{ij}=G-\{v_{1},v_{2},v_{3}\}-N[y_{i},v_{j}]$, we have $\operatorname{minsurp}(G^{\prime}_{ij})\geq-3+\operatorname{codeg}(y_{i},v_{j})$. Moreover, if $\operatorname{codeg}(y_{i},v_{j})=1$ and $\operatorname{minsurp}(G^{\prime}_{ij})=-2$, then either $S(G^{\prime}_{ij})\geq 1$ or $S(v_{j^{\prime}},N[v_{j}])\geq 1$ for some $j^{\prime}\in\{1,2,3\}$.
@@
Now suppose $\operatorname{codeg}(y_{1},v_{1})=1$ and $\operatorname{surp}_{G^{\prime}}(I)=-2$. If $|I|\geq 3$ then $S(G^{\prime})\geq 1$ as desired, so suppose $|I|=2$. By Observation 62, we have $\operatorname{codeg}(v_{2},v_{3})\leq 2$, so there must be some vertex $x\in I$ and $v_{j}\in\{v_{2},v_{3}\}$ with $x\not\sim v_{j}$; say without loss of generality $x\not\sim v_{3}$ and $x$ is isolated in $G^{\prime}$.
Let $d=\deg_{G}(x)$. By Observation 62 we have $\operatorname{codeg}(x,y_{1})\leq d-3$ and $\operatorname{codeg}(x,v_{1})\leq 2$. So the only way for $x$ to be isolated in $G^{\prime}$ is to have $x\sim v_{2},\operatorname{codeg}(x,y_{1})=d-3,\operatorname{codeg}(x,v_{1})=2$. If $d\leq 5$, we thus have $S(v_{2},N[v_{1}])\geq 1$ due to subquadratic vertex $x$. If $d=6$, then by Proposition 66 we have $x\not\sim y_{2}$ (else the 6-vertex $x$ would be shared by $y_{2},v_{2}$); in this case, GN[x]G-N[x] has a 3-vertex $y_{1}$ with a triangle $y_{1},y_{2},u$, and we would branch on $x$ at Line 2. ∎
@@
Case I: $\boldsymbol{y_{1}\not\sim y_{2}}$. Split on $y_{1},y_{2},v_{1}$, getting subproblems $G_{1}=G-N[y_{1}],G_{2}=G-y_{1}-N[y_{2}],G_{3}=G-\{y_{1},y_{2}\}-N[v_{1}],G_{4}=G-\{y_{1},y_{2},v_{1}\}$. By Observation 62, we have $\operatorname{minsurp}(G_{1})\geq 2$ and $\operatorname{minsurp}(G_{2})\geq 1$ and $\operatorname{minsurp}(G_{3})\geq\operatorname{shad}(y_{1},v_{1})-1\geq 1$ and $\operatorname{minsurp}(G_{4})\geq\operatorname{shad}(y_{1},y_{2})-1\geq 1$. So $G_{1},G_{2},G_{3},G_{4}$ have drops $(2.5,6),(3,7),(3,7),(1.5,3)$ directly. We also apply Proposition 30 to 2-vertex $u$ in $G_{4}$, where $\deg(v_{2})+\deg(v_{3})-1-\operatorname{codeg}(v_{2},v_{3})\geq 7$ by Observation 62. So we need:
@@
Case II: $\boldsymbol{y_{1}\sim y_{2}}$. We choose some enumeration $v_{1},v_{2},v_{3}$, and split on $v_{1},v_{2}$; in the subproblem $G-\{v_{1},v_{2}\}$, we apply (B-Fun2) to $u,y_{1}$. This generates subproblems $G_{1}=G-N[v_{1}],G_{2}=G-v_{1}-N[v_{2}],G_{3}=G\{v_{1},v_{2}\}-N[v_{3},y_{1}],G_{4}=G-\{v_{1},v_{2},y_{1}\}$.
By Observation 62, $G_{1},G_{2},G_{4}$ have drops $(2,5),(2.5,6),(1.5,3)$ directly; furthermore, we can apply Proposition 30 to the 2-vertex $u$ in $G_{4}$, where $r=\deg(v_{1})+(\deg(y_{2})-1)-\operatorname{codeg}(y_{2},v_{1})-1\geq 5+5-2-1=7$ by Observation 62.
Proposition 67 and Observation 62 give $\operatorname{minsurp}(G_{3})\geq-3+\operatorname{codeg}(y_{1},v_{3})$ and $\operatorname{codeg}(y_{1},v_{3})\leq 2$. In light of Proposition 66, there are two cases to select the ordering:
Case IIa: $\boldsymbol{S(v_{1},N[v_{2}])\geq 1.}$ In this case $G_{2}$ has drop $(2.5,7)$ after simplification. Since $G_{3}$ has drop $(4,11)$, we need:
@@
Case IIb: $\boldsymbol{\operatorname{codeg}(y_{1},v_{3})=1.}$ By Proposition 67, either $S(v_{j},N[v_{3}])\geq 1$ for some $j$ (which would put us into Case IIa, with some other vertex ordering), or $S(G_{3})\geq 1$, or $\operatorname{minsurp}(G_{3})\geq-1$. In the latter two cases, $G_{3}$ has drop $(4,13)$ or $(4.5,12)$. So we need:
@@
1Simplify $G$
2Branch on a vertex $u$ with degree larger than 7
3Split on a 7-vertex $u$ with $\operatorname{shad}(N[u])\geq 0$.
4Branch on a 7-vertex $u$ which has a blocker $x$ with $\deg(x)\geq 5$.
5Branch on a vertex $x$ which is a blocker of three 7-vertices $u_{1},u_{2},u_{3}$.
6Run either Branch6-3 or the MaxIS-7 algorithm.
Algorithm 10 Function Branch7(G,k)(G,k)
@@
Line 2. Here, we apply Figure 6; we need:
@@
Line 3. In this case GN[u]G-N[u] has drop $(3,7)$. So we need:
@@
Line 4. Apply (B-Block) to $u,x$, getting subproblems $G-N[x],G-\{u,x\}$ with drops $(1,5)$ and $(1,2)$ respectively. We need:
@@
Line 5. Apply (B-Block) to $u_{1},u_{2},u_{3},x$. The resulting subproblems $G-\{u_{1},u_{2},u_{3},x\}$ and GN[x]G-N[x] have drops $(1,4),(1,3)$ directly; the branch-seq $[(1,3),(1,4)]$ is already covered in (7-1).
@@
1If $\operatorname{maxdeg}(G)\leq 7$, then use Lemma 15 with the algorithms Branch7 or MaxIS-7.
Otherwise, split on a vertex of degree at least $8$.
Algorithm 11 Function Branch8(G,k)(G,k)
@@
Splitting on a vertex of degree $r\geq 8$ generates subproblems with vertex covers of size at most k1k-1 and k8k-8. These have cost $1.25284^{k-1}$ and $1.25284^{k-8}$ by induction. Since $1.25284^{-1}+1.25284^{-8}<1$, the overall cost is $O^{*}(1.25284^{k})$. Otherwise, Lemma 15 (with parameters $a_{7},b_{7}$ and $c=\log(1.19698)$) gives the desired algorithm with runtime $O^{*}(1.25284^{k})$. ∎
@@ Appendix A Proofs for basic results
Proof of Proposition 4 and Proposition 5.
Consider the linear program defined by $\operatorname{LPVC}(G)$ and the additional constraints $\theta(v)=0:v\in I$ and $\theta(v)\geq 1/2:v\notin X$, and let its optimum value be denoted by $\lambda_{I,X}$. We claim that any indset $J$ of minimum surplus with $I\subseteq J\subseteq X$ corresponds to an optimal basic solution to this linear program, and has $\lambda_{I,X}=\tfrac{1}{2}(n+\operatorname{surp}_{G}(J))$. This will show Proposition 5 (since the linear program can be solved in polynomial time), and it will also show Proposition 4 (since $\lambda(G)=\lambda_{\emptyset,V}=\operatorname{minsurp}^{0}(G)$).
First, we claim that any such basic solution $\theta\in[0,1]^{V}$ is half-integral. For, suppose not, and let $A,B$ denote the vertices $v$ with $\theta(v)\in(0,1/2)$ and $\theta(v)\in(1/2,1)$ respectively. It can be seen that modifying $\theta(v):v\in A$ by $\pm\epsilon$ and $\theta(v):v\in B$ by $\mp\epsilon$ preserves feasibility. Thus, in a basic solution, we must have $A=B=\emptyset$.
Now consider any such basic solution $\theta\in\{0,\tfrac{1}{2},1\}^{V}$. Let $I_{0}=\theta^{-1}(0)$ and $I_{1}=\theta^{-1}(1)$; note that $I_{0}$ is an indset (possibly empty) and $N(I_{0})\subseteq I_{1}$. So $\sum_{v}\theta(v)=|I_{1}|+\tfrac{1}{2}(n-|I_{0}|-|I_{1}|)\leq(n+|N(I_{0})|-|I_{0}|)/2$. On the other hand, for any min-set $J$ with $I\subseteq J\subseteq X$, then setting $\theta(v)=1$ for $v\in N(J)$ and $\theta(v)=0$ for $v\in J$ and $\theta(v)=1/2$ for all other vertices $v$ gives a fractional solution cover with $\sum_{v}\theta(v)=(n+\operatorname{surp}_{G}(J))/2$. ∎
@@
Let $I$ be a critical-set; we first claim there is a good cover $C$ with $I\subseteq C$ or $I\cap C=\emptyset$. For, suppose that $I\cap C=J$ where $\emptyset\neq J\subsetneq I$. In this case, we must also have $N(I\setminus J)\subseteq C$, so overall $|C\cap N[I]|\geq|J|+|N(I\setminus J)|=|J|+|I\setminus J|+\operatorname{surp}(I\setminus J)$; since $I\setminus J$ is a non-empty subset of II, we have $\operatorname{surp}(I\setminus J)\geq\operatorname{surp}(I)$, hence this is at least $|J|+|I\setminus J|+\operatorname{surp}(I)=|I|+\operatorname{surp}(I)\geq N(I)$. Thus, replacing $C\cap N[I]$ by $N(I)$ would give a vertex cover $C^{\prime}$ with $C^{\prime}\cap I=\emptyset$ and $|C^{\prime}|\leq|C|$.
@@
Proof of Observation 9.
Here $\operatorname{minsurp}(G)\geq 2$. If $J$ is a min-set of GXG-X, then $2\leq\operatorname{surp}_{G}(J)\leq\operatorname{surp}_{G-X}(J)+|X|$. So $\operatorname{surp}_{G-X}(J)\geq 2-|X|$. Similarly, if $J$ is a min-set of GN[I]G-N[I], then $2\leq\operatorname{surp}_{G}(I\cup J)=\operatorname{surp}_{G-N[I]}(J)+\operatorname{surp}_{G}(I)$. ∎
Proof of Proposition 10.
Proof of Proposition 11.